OpenStudy (anonymous):

Using a directrix of y = -2 and a focus of (1, 6), what quadratic function is created?

OpenStudy (anonymous):

A.f(x) = one eighth (x - 1)2 - 2 B. f(x) = -one eighth (x + 1)2 - 2 C. f(x) = -one sixteenth (x + 1)2 - 2 D. f(x) = one sixteenth (x - 1)2 + 2

OpenStudy (anonymous):

I know it!

OpenStudy (anonymous):

Know what? lol

OpenStudy (anonymous):

@Brostep0s

OpenStudy (triciaal):

was the answer D?

OpenStudy (anonymous):

I don't know I reposted the question no one helped me

OpenStudy (anonymous):

Yes

OpenStudy (triciaal):

I was not sure so won't post unless reviewed and sure